OpenStudy (anonymous):

X^4=2x^2-1 solve algebracally

myininaya (myininaya):

\[x^4-2x^2+1=0\] \[(x^2)^2-2x^2+1=0\] replace x^2 with u and see if you can figure out now let me know if you still have trouble after replacing with u

OpenStudy (anonymous):

well then if you use the quadratic form its not real is that correcr?

myininaya (myininaya):

\[u^2-2u+1=0\] this is factorable right? \[(u-1)(u-1)=0\]

myininaya (myininaya):

u-1=0 =>u=1 but u=x^2 x^2=1=>x=1 or x=-1

OpenStudy (anonymous):

so one more thing to ask for factoring what number am I factoring

myininaya (myininaya):

you are factoring the polynomial u^2-2u+1 you can use quadratic and will give you one real solution

myininaya (myininaya):

quadratic formula*

OpenStudy (anonymous):

yeah yeah, but where did you get the 1 and 1 from (u−1)(u−1)=0

myininaya (myininaya):

\[u=\frac{2 \pm \sqrt{(-2)^2-4(1)(1)}}{2(1)}=\frac{2 \pm \sqrt{4-4}}{2}=\frac{2 \pm 0}{2}=\frac{2}{2}=1\]

OpenStudy (anonymous):

ok got it, thanks

myininaya (myininaya):

I used the following process to factor: u^2+bu+c only works when coefficient of u^2 is 1 find factors of c that add up to be b then put the numbers in the blanks below (u+__)(u+__)

OpenStudy (anonymous):

u=2±(−2)2−4(1)(1)−−−−−−−−−−−−√2(1)=2±4−4−−−−√2=2±02=22=1

OpenStudy (anonymous):

at the second part how did you get I got -8 under the radical how did you get 0

myininaya (myininaya):

(-2)^2=(-2)(-2)=4 4-4=0

myininaya (myininaya):

\[u=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

oh ok never mind

myininaya (myininaya):

i think you forgot to square the negative part thats why you got -8